bc(b-c)yzp+ca(c-a)xzq=ab(a-b)xy


1
Expert's answer
2021-01-12T14:49:07-0500

P=bc(b-c)yz=b2cyz - bc2yz

Q=ca(c-a)xz=c2axz - ca2xz

R=ab(a-b)xy=a2bxy - ab2xy

"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"

"\frac{dx}{b^2cyz - bc^2yz}=\frac{dy}{c^2axz - ca^2xz}=\frac{dz}{a^2bxy - ab^2xy}"

Multiplyers: ax,by,cz

axdx+bydy+czdz=0

1/2(ax2+by2+cz2)=C1

Multiplyers: -ax,-by,-cz

-axdx-bydy-czdz=0

-1/2(ax2+by2+cz2)=C2

The general solution is,

"\phi(c_{1},c_{2})=0"

"\phi( 1/2(ax^2+by^2+cz^2),-1/2(ax^2+by^2+cz^2))=0"